Disclaimer: This material is being kept online for historical purposes. Though accurate at the time of publication, it is no longer being updated. The page may contain broken links or outdated information, and parts may not function in current web browsers. Visit https://espo.nasa.gov for information about our current projects.



3/1 Accomplishments:
AV-6 landing gear retraction and a gain in altitude
A/C currently flying @ 4°N, 100°W, approximately 13 hours into ATTREX's last science flight

3/2 Plans:
Recover the a/c and begin instrument removal


2/28 Accomplishments:
Ku system problems were unresolved but the decision was made to proceed with last ATTREX science flight
Instruments re-installed
SSFR calibration complete
A/C preflight underway

3/1 Plans:
Get the aircraft off the ground
Conduct a successful final science flight, destination is the 2.5N, 85W area.

Investigator hands-off @ 02:00, a/c roll-out @ 03:00, engine start @ 05:00, take-off @ 0600.

Instrument teams must have a representative in the GHOC @ 04:45.


2/27 Accomplishments:
A/C landed at approximately 10:20 local time after a 24.4 hour flight
During the flight the Ku system problem was traced to either a High Power Amplifier (HPA) problem or a wave guide problem
NOAA water vapor and O3, UCATS instruments were removed
UCATS experienced an anomaly in their methane and hydrogen signals that they're investigating
GWAS changed cans and has completed their preflight

2/28 Plans:
Roll the a/c out at 06:00 to begin Ku system tests
Re-install instruments
SSFR calibration
Instrument and a/c preflights

The plan is to take-off @ 06:00 on Friday morning for our last science flight.  That means aircraft will roll out @ 03:00.  Investigator hands off will be at 02:00 Friday morning!!!!!!!


2/26 Accomplishments:
Despite MPCS modem problems that put us on the verge of canceling the flight the aircraft took off three hours late on our 5th science flight
We experienced Ku satellite communication problems and so had no Ku for the flight (problems on the aircraft are the initial thought)
Due to ambient temperatures along the route that were higher than predicted the aircraft is burning more fuel than expected and will have to turn for home sooner than planned
Despite these issues the aircraft will get to approximately 7°N, 170°W before executing a U turn
All instruments appear to be working well

2/27 Plans:
Recover, weigh and refuel the aircraft
Remove those instruments that require removal
Begin to investigate the Ku problem

The plan is to conduct our last science flight on Friday/Saturday


2/25 Accomplishments:
All instrument problems fixed
All instruments preflighted
A/C preflighted

2/25 Plans:
07:00 Take-off for our 5th science flight (estimated duration of 25 hours)


2/22 Accomplishments:
A/C landed at 07:24 after a 24.6 hour flight
NOAA O3 and water vapor removed (O3 has a computer problem)
NOAA UCATS removed
mini-DOAS removed (to work a problem with the IR telescope)

2/23 Plans:
Work on instrument problems
Field trip of 99 students from the Palmdale Aerospace Academy Middle School

Next week's preliminary plan to fly on Monday evening, with a 20:00 PST take-off, has been changed to flying on Tuesday with an 07:00 PST take-off.  The next, and last, flight would then be scheduled for Friday-Saturday.


2/21 Accomplishments:
A/C took off 13 minutes early for ATTREX science flight #4
A/C currently on the return leg of the science flight that took it to 12S latitude

2/22 Plans:
Recover A/C
Weigh and fuel the A/C
Prepare for Saturday field trip of 99 students from the Palmdale Aerospace Academy Middle School

Science Flight #5 is planned for Monday with an ETD of 20:00 PST


2/20 Accomplishments:
Instrument preflights completed
A/C preflights completed

2/21 Plans:
03:00 Science Team hands off aircraft
05:45 Staff GHOC
07:00 take-off for ATTREX science flight #4

Instrument teams need to have a representative in the GHOC @ 05:45!


2/8 Accomplishments:
GWAS reinstalled and preflighted
NOAA UCATS reinstalled and preflighted
NOAA O3 lamp changed (I miss-stated yesterday that the lamp change was in UCATS) and it and Water Vapor are reinstalled and preflighted.
All other instruments preflighted
Aircraft preflighted

2/9 Plans:
06:30 take-off for second science flight

Instrument teams need to have a representative in the GHOC @ 05:15!!


2/7 Accomplishments:
GWAS determined their pump problem was a frozen bearing.  They think it was because the pump got too hot so they're going to use thermistors to monitor the pump's temperature.
While addressing a noisy pressure sensor NOAA UCATS got indications of a failing mercury lamp.  Pressure sensor was changed, lamp is on the way to be changed tomorrow afternoon
Got approval from all of those required to attempt a Saturday/Sunday science flight.
HU/PCRS removed their gas deck, beefed up a solenoid driver to ensure robust calibrations and reinstalled the unit.
Mini-DOAS had their telescope fairing reworked to address interference issues
SSFR conducted a system vacuum check and calibration
MMS submitted data to the ATTREX archive

2/8 Plans:
Reinstall GWAS pump pallet
Reinstall NOAA O3 and  Water Vapor and UCATS
Perform instrument and aircraft preflights


Subscribe to Daily Schedule